Oxygen (molar mass 32 kg/kmol) expands reversibly in a cylinder behind a piston at a constant pressure of 3 bar. The volume initially is 0.0114 mand finally is 0,0319m3 , the initial temperature is 16.380C. Calculate the absolute temperature after expansion with the correct unit. Assume oxygen to be a perfect gas and take the specific heat at constant pressure as = 0.917kj/kgK.


1
Expert's answer
2021-08-17T08:24:01-0400

molar mass = 32 kg/kmol

P = constant

P = 3 bar

V1 = 0.01m3                  V2 = 0.03m3

TI= 170C  = 290.15K

W = ?

Q = ?

Constant Pressure (Cp) = 0.917kj/kgK

 

Therefore,

W = P(V2 – V1) = 3 × 102 (0.0319 – 0.0114)

W = 6.15 KJ/6150J

3⋅105 (0.01) = m/32 (8.31) (290)

m = 39.8165g/ 0.039Kg

n = 39.8165g/32

n = 1.2443 mol

Q = m Cp ∆T

Q = 0.039 × 0.917 × (870 – 290.15)

Q = 20. 74KJ/ 20740J

 

3.105 (0.03) = m/32 (8.31) T2

T2 = 870k